OpenStudy (rahulmr):

Differentiate \[\frac{ x^2 }{ 4 }\log_{e}x \] and hence find \[\int\limits x \log_{e}x \]. Help please. I don't know to should i do this question. Can anyone help me out. Thanks.

OpenStudy (sshayer):

\[\frac{ d }{ dx }\left( \frac{ x^2 }{ 4 }\log_{e}x \right)=\frac{ x^2 }{ 4 }*\frac{ 1 }{ x }+\frac{ 2x }{ 4 }\log_{e } x\] \[=\frac{ x }{ 4 }+\frac{ 2x }{ 4 }\log_{e} x\]

OpenStudy (rahulmr):

I found how to do it. Thanks for the help.

OpenStudy (rahulmr):

the answer was \[ \frac{ x^2\log_{e}x }{ 2 } -\frac{ x^2 }{ 4 }\]

OpenStudy (sshayer):

\[\int\limits \left( \frac{ x }{ 4 }+\frac{ 2x }{ 4 }\log_{e}x \right)dx=\frac{ x^2 }{ 4 }\log_{e} x+c\] now you can complete it.

OpenStudy (rahulmr):

Thanks.

OpenStudy (sshayer):

yw.